Please show me the steps to solve this for the vertex, axis of symmetry, focus and directrix:
y-4=-1/16(x+1)^2
Answer by lwsshak3(11628) 
 solve this for the vertex, axis of symmetry, focus and directrix:
y-4=-1/16(x+1)^2
(x+1)^2=-16(y-4)
This is an equation of a parabola that opens down.
Its basic form of equation: (x-h)^2=4p(y-k), (h,k)=coordinates of the vertex
For given parabola:
vertex: (-1,4)
axis of symmetry: x=-1
4p=16
p=4
focus: (-1,0) (p-distance below vertex on the axis of symmetry)
directrix: y=8 (p-distance above vertex on the axis of symmetry)
|
|
|